As of 8pm local time, the IFO seems to have equilibriated to atmospheric pressure (I don't hear the hiss of in-rushing air near 1X8 and P1a reports 760 torr). The pumpspool looks healthy and there are no signs in the TP diagnostics channels that anything bad happened to the pumps. Chub is working on getting the N2 setup more robust, we plan to take the EY door off at 9am tomorrow morning with Bob's help.
* I took this opportunity to follow instructions on pg 29 of the manual and set the calibration for the SuperBee pirani gauge to 760 torr so that it is in better agreement with our existing P1a Pirani gauge. The correction was ~8% (820-->760).